# Video Deepfake Abuse: How Company Choices Predictably Shape Misuse Patterns

### v1

@@ 5.2. Policy Influences Company Incentives
This paper has focused on how the choices made by developers and model distribution platforms will affect downstream misuse patterns for AI video generators. However, it is key, especially for policy audiences, to recognize that these choices are not made in a vacuum, but amidst external commercial, regulatory, and liability incentives. Diffusion model generators and distribution platforms have commercial incentives to deploy and share models as quickly as possible, but regulatory requirements, liability, and transparency can have a significant counteractive effect to incentivize risk mitigation (bommasani_considerations_2024; williams2025regulating; smith2024liability). When considering regulation related to managing AI risks– especially from open-weight models, policymakers should recognize that some types of risk management techniques, such as staged deployment, may be more onerous for small developers and academics than for larger companies.
Some laws on AIG-NCII have been written to regulate content and models. Some regulatory approaches have aimed to influence the AIG-NCII ecosystem by targeting downstream non-consensually distributed AIG-NCII, such as the federal TAKE IT DOWN Act in the USA (sen_cruz_ted_r-tx_take_2023). Other approaches can also focus on models themselves. For example, Arkansas Act 827, enacted April 2025, states that “the Attorney General may institute a civil action on behalf of the state against a provider or developer of image generation technology that was used to create deepfake visual material” if “the provider or developer of the image generation technology did not have reasonable safeguards in place to protect against the generation of deepfake visual material” (state_of_arkansas_act_2025).
Transparency requirements invite external scrutiny. Another key factor in the incentives of developers and model distributors is transparency (widder2022limits). Detailed information about risk management measures gives external stakeholders the ability to study their mitigations. A low-transparency state of affairs, as illustrated in Table 1, can be expected by default because selective transparency helps companies avoid unwanted forms of scrutiny (ananny2018seeing). For example, the lack of information about AIG-NCII risk mitigations from developers (see Table 1) impedes work like ours from directly identifying specific links between company choices and downstream harms. Absent external incentives, a low transparency regime will tend to self-reinforce by inviting increased levels of scrutiny and criticism to developers who voluntarily choose to be more transparent.
@@ 5.3. Weighing Competing Considerations
Balancing the benefits and risks of open-weight models: As we have discussed, models with openly available parameters are a key factor and key driver behind the creation of AIG-NCII and AIG-CSAM (iwf2023ai-csam; widder2022limits). However, this does not imply open-weight models are net harmful. They are also key for positive uses, including the de-concentration of power and the facilitation of open research on safety (kapoor_societal_2024; bommasani_considerations_2024; francois_different_2025). More pragmatically, powerful open-weight models are increasingly influential (bhandari2025forecasting) and likely inevitable. In recognition of this, we have emphasized a harm mitigation approach for safeguarding open-weight models in Section 4, arguing that substantial reduction of AIG-NCII harms can be achieved without prohibiting powerful open-weight models.
Balancing the benefits and risks of NSFW AI Content: AIG-NCII is a serious form of image-based sexual abuse that results in psychological, financial, and reputational harm, disproportionately borne by women and girls (american_sunlight_project_deepfake_2024; henry_ajder_state_2019; flynn_deepfakes_2022). Prior works have argued that the current NSFW generative AI ecosystem is characterized by non-consensual practices (gibson2025analyzing; ding2024malicious). However, AI models that produce NSFW content could also have applications based on consensual and ethical use. (geeng2025co). Work to build infrastructure for more positive uses of NSFW AI content between adults would require consent-focused data collection, conscientious deployment, and concerted effort to reduce AIG-NCII and AIG-CSAM risks (geeng2025co; Cintaqia2025Stop).
@@ Appendix A Methodological Details
Img2Vid and Framepacking tools: in Section 3.2, we restrict our analysis to AI video generators. However, it is important to note that a substantial fraction of AI videos online come from image generators scaffolded using tools such as Img2Vid and Framepack (zhang2025packing). However, these techniques are generally more time-consuming and typically produce lower-quality video than modern video generators.

### v2

@@ 5.2. Policy Influences Company Incentives
This paper has focused on how the choices made by developers and model distribution platforms will affect downstream misuse patterns for AI video generators. However, it is key, especially for policy audiences, to recognize that these choices are not made in a vacuum, but amidst external commercial, regulatory, and liability incentives. Diffusion model generators and distribution platforms have commercial incentives to deploy and share models as quickly as possible, but regulatory requirements, liability, and transparency can have a significant counteractive effect to incentivize risk mitigation (Bommasani et al., 2024; Smith et al., 2024; Ramakrishnan et al., 2024; Williams et al., 2025). When considering regulation related to managing AI risks– especially from open-weight models, policymakers should recognize that some types of risk management techniques, such as staged deployment, may be more onerous for small developers and academics than for larger companies.
Some laws on AIG-NCII have been written to regulate content and models. Some regulatory approaches have aimed to influence the AIG-NCII ecosystem by targeting downstream non-consensually distributed AIG-NCII, such as the federal TAKE IT DOWN Act in the USA (Sen. Cruz, Ted [R-TX], 2023). Other approaches can also focus on models themselves. For example, Arkansas Act 827, enacted April 2025, states that “the Attorney General may institute a civil action on behalf of the state against a provider or developer of image generation technology that was used to create deepfake visual material” if “the provider or developer of the image generation technology did not have reasonable safeguards in place to protect against the generation of deepfake visual material” (State of Arkansas, 2025).
Transparency requirements invite external scrutiny. Another key factor in the incentives of developers and model distributors is transparency (Widder et al., 2022). Detailed information about risk management measures gives external stakeholders the ability to study their mitigations. A low-transparency state of affairs, as illustrated in Table 1, can be expected by default because selective transparency helps companies avoid unwanted forms of scrutiny (Ananny and Crawford, 2018). For example, the lack of information about AIG-NCII risk mitigations from developers (see Table 1) impedes work like ours from directly identifying specific links between company choices and downstream harms. Absent external incentives, a low transparency regime will tend to self-reinforce by inviting increased levels of scrutiny and criticism to developers who voluntarily choose to be more transparent.
@@ 5.3. Weighing Competing Considerations
Balancing the benefits and risks of open-weight models: As we have discussed, models with openly available parameters are a key factor and key driver behind the creation of AIG-NCII and AIG-CSAM (Internet Watch Foundation (IWF), 2023; Widder et al., 2022). However, this does not imply open-weight models are net harmful. They are also key for positive uses, including the de-concentration of power and the facilitation of open research on safety (Kapoor et al., 2024; Bommasani et al., 2024; François et al., 2025). More pragmatically, powerful open-weight models are increasingly influential (Bhandari et al., 2025) and likely inevitable. In recognition of this, we have emphasized a harm mitigation approach for safeguarding open-weight models in Section 4, arguing that substantial reduction of AIG-NCII harms can be achieved without prohibiting powerful open-weight models.
Balancing the benefits and risks of NSFW AI Content: AIG-NCII is a serious form of image-based sexual abuse that results in psychological, financial, and reputational harm, disproportionately borne by women and girls (American Sunlight Project, 2024; Henry Ajder et al., 2019; Flynn et al., 2022). Prior works have argued that the current NSFW generative AI ecosystem is characterized by non-consensual practices (Gibson et al., 2025; Ding and Suresh, 2025). However, AI models that produce NSFW content could also have applications based on consensual and ethical use. (Geeng et al., 2025). Work to build infrastructure for more positive uses of NSFW AI content between adults would require consent-focused data collection, conscientious deployment, and concerted effort to reduce AIG-NCII and AIG-CSAM risks (Geeng et al., 2025; Cintaqia et al., 2025).
@@ 5.4. Limitations
Ethical Considerations
Our research involved analyzing publicly accessible content, including NSFW material. We adhered to the following ethical guidelines: (1) We did not seek, download, store, or view non-consensual intimate imagery, and (2) while we name platforms and models to enable policy action, we have withheld specific search terms, links, and website names that could drive traffic toward websites that may facilitate harm.
@@ Appendix A Methodological Details
Img2Vid and Framepacking tools: in Section 3.2, we restrict our analysis to AI video generators. However, it is important to note that a substantial fraction of AI videos online come from image generators scaffolded using tools such as Img2Vid and Framepack (Zhang and Agrawala, 2025). However, these techniques are generally more time-consuming and typically produce lower-quality video than modern video generators.